OpenStudy (angela210793):

how do you solve X^3-4x+3=0

OpenStudy (rational):

by inspection, 1 is a root

hartnn (hartnn):

will you be able to solve if you know one of the roots ?

hartnn (hartnn):

here its very easy to find, SUM of co-efficients = 0 hence x=1 must be the root

OpenStudy (rational):

synthetic division and quadratic formula are your best friends here

hartnn (hartnn):

agreed, synthetic or long division i'd go for synthetic

OpenStudy (imtiaz7):

yes synthetic division it's the best method to solve this

hartnn (hartnn):

for ANY polynomial, its easy to determine whether x=1 or x=-1 is one of the solution if sum of ALL co-efficients = 0 , then x= 1 is a root if sum of co-efficients of even exponents of x = sum of co-efficient of odd exponents of x, then x=-1 is the root

OpenStudy (angela210793):

I know how to solve quadratic eq...i just dont get how did u know 1 was a root?? O.o did u just guess? O.o

hartnn (hartnn):

no guess this is acceptable reason if sum of ALL co-efficients = 0 , then x= 1 is a root

OpenStudy (rational):

+1 and -1 are the first things to guess and check

hartnn (hartnn):

if you still don't want to use that you can use rational root theorem

OpenStudy (rational):

a cubic has a direct formula but its very messy and very very hard to solve it analytically. so you should always try +1, -1 and every other rational root thats predicted by rational root theorem

OpenStudy (angela210793):

woooow...i didnt know that O.o Thanks a lot Hartnn :D

hartnn (hartnn):

if you know it, you will need to check for just these values +1,-1,+3,-3

OpenStudy (rational):

also descartes rule of signs makes ur life a bit easy by eliminating few of the roots (doesn't matter for current problem as 3 is a prime)

hartnn (hartnn):

you're welcome ^_^ you'll get a quadratic after synthetic division also, you can have a look here http://openstudy.com/study#/updates/504335bee4b000724d4620b4

hartnn (hartnn):

for \(ax^3+bx^2+cx+d =0\) a,b,c,d are the co-efficients if a+b+c+d = 0, then x= 1 is the root example : x^3 - 4x+6x-3 what are the co-efficients ?

OpenStudy (angela210793):

1 -4 and 3 = 0 thankssss :")

hartnn (hartnn):

1, -4 , 6 , -3 add them, you get 0

hartnn (hartnn):

similarly, let me give you an example for x=-1 x^3 - 3x^2 +3x - 1 what are co -efficients of x^3 and x ?? add them what are co -efficients of x^2 and constant ?? add them

OpenStudy (angela210793):

1-3+3-1 = 0 so whenever it goes equal to 0, it means the root is 1? what about the other thing that x-1 is the factor?

hartnn (hartnn):

ok, i messed up with the example for this case too, x=1 is one of the solution, yes because sum of co-efficients is 0 now if 'a' is a root, (x-a) is a factor!

OpenStudy (angela210793):

no prob lol :P i understood the main part :) it's a factor so i can factorize the main equation?

hartnn (hartnn):

yes i mean you can use that for synthetic division you'll use x=1

hartnn (hartnn):



OpenStudy (angela210793):

and where did those 3 come from? O.o ://

hartnn (hartnn):